OpenStudy (anonymous):

Given F = 9 5 C + 32, the conversion formula for Fahrenheit to Celsius, solve for C.

OpenStudy (anonymous):

Jeffrey is training to run a marathon. On Monday he ran and walked a total of 11 miles. If he ran 2 miles more than twice as far as he walked, how many miles did he run?

OpenStudy (anonymous):

You just have to re arrange that equation, to make C equal to the rest of the equation. F = 95 c + 32 F - 32 = 95c F - 32 / (95) = c

OpenStudy (anonymous):

Jeffrey is training to run a marathon. On Monday he ran and walked a total of 11 miles. If he ran 2 miles more than twice as far as he walked, how many miles did he run?

OpenStudy (anonymous):

easier if you set up a equation r = running , w = walking r + w =11 r = 2 + 2w solve for w 2 + 2w + w = 11 3w = 11 - 2 w = 9 / 3 w = 3 miles this means he walked 3 miles if we know how far he walked we can figure out how far he ran uysing our original equation w + r =11 3 + r = 11 r = 11 - 3 r = 8 miles